E6-4 (L03,4) (Computation of Future Values and Present Values) Using the appropriate interest table, answer the following questions. (Each case is independent of the others).

  1. (a) What is the future value of 20 periodic payments of $4,000 each made at the beginning of each period and compounded at 8%?

  2. (b) Whatisthepresentvalueof$2,500tobereceivedatthebeginningofeachof30periods,discountedat5%compoundinterest?

  3. (c) What is the future value of 15 deposits of $2,000 each made at the beginning of each period and compounded at 10%?

    (Future value as of the end of the fifteenth period.)

  4. (d) Whatisthepresentvalueofsixreceiptsof$1,000eachreceivedatthebeginningofeachperiod,discountedat9%com-

    pounded interest?
Answer #1

Using excel PV and FV function as follows:

a)

=FV(8%,20,4000,,1)

=197691.69

b)

=PV(5%,30,2500,,1)

=40352.68

c)

=FV(10%,15,2000,,1)

=69899.46

d)

=PV(9%,6,1000,,1)

=4889.65